Translate: Eight times the difference between four times a number and five is equal to twenty more than twice the number. Not sure if my translation is correct. 8(4n)+5=20+2n
Found 2 solutions by fractalier, AnlytcPhil:
Answer by fractalier(2101)

You are close...you should have written
8(4n - 5) = 20 + 2n

Answer by AnlytcPhil(1116)

Translate: Eight times the difference between four times a number and
five is equal to twenty more than twice the number. Not sure if my 
translation is correct.   8(4n)+5=20+2n

>>...Eight times the difference between four times a number and five 
is equal to twenty more than twice the number...<<

Replace the words "a number" and "the number" by n, so we now have:

>>...Eight times the difference between four times n and five is equal
to twenty more than twice n...<<

>>...Eight times the difference between four times n and five is equal
to twenty more than twice n...<<

Replace the words "four times n" by "4n" and "twice n" by "2n", so we 
now have:

>>...Eight times the difference between 4n and five is equal
to twenty more than 2n...<<

>>...Eight times the difference between 4n and five is equal
to twenty more than 2n...<<

Replace the words "the difference between 4n and five" by "4n-5" and 
"twenty more than 2n" by "2n+20", so we now have:

>>...Eight times 4n-5 is equal to 2n+20...<<

>>...Eight times 4n-5 is equal to 2n+20...<<

Replace the words "Eight times 4n-5" by "8(4n-5)" and 
"is equal to" by "=", so we now have:

>>...8(4n-5) = 2n+20...<<

     8(4n-5) = 2n+20

Solve that and get

n = 2

Edwin